BGA has 345,000 shares outstanding that sell for $30 per share. The company plans a 3-for-5 reverse stock split. How many shares will be outstanding after the split?
Answer #1

Initial Number of Shares =345000
Shares outstanding after that split =Number of Shares *3/5=345000*3/5 =207000